BI HiRadMat
→
Europe/Zurich
Description
Join Zoom Meeting
https://cern.zoom.us/j/62435854052?pwd=TE5mVzR4VVJxL1RSZWs4Mkl2eHNFUT09
Hi,
Agenda for this meeting:
- BI Electronics infrastructure status – (Stephane – 5min)
- BPM experiment status (?? – 5min)
- BTV reinstallation status – (Stephane – 5min)
- Nano tube experiment status (Alexandre – 5 min)
- Pulse lists discussion
- Readiness of the 3 exp. for the HRM TB of the 25th of May (safety file, etc…)
- Reminder: Safety files should be sent to the TB (via Pascal) max 1 week before the FB
- BTV: in principle no change
- CHEFT: Updated draft already sent by Alexandre
- BPM: should be updated with mechanical integration
- Reminder: Safety files should be sent to the TB (via Pascal) max 1 week before the FB
- AOB
Cheers,
Stephane
- List of cables not available for the moment. We can contact Pascal (and Aymeric to check form time to time for the moment). I hope it will be available soon.
- A request to replace the 3U VME with a 9U VME was done
- Confirmation of Pascal "The three BI experiments have been allocated a total of 3 beamtime shifts (3*8 hours). How you share this in-between you guys is up to you". We can make it clear internally next meeting.
- Cherenkov pick-ups: Still to be defined if the acquisition is to be done in TT61 or in surface
There are minutes attached to this event.
Show them.